OpenStudy (anonymous):

I just need someone to check if these are correct? Please & Thanks. 1. Use the substitution method to solve the system of equations. x + 2y = 15 2x – 3y = 2 a. (2, 4) b. (4, 7) c. (7, 4) <---- i chose d. (4, 2) 2. When using the substitution method, what expression for y in the second equation would be substituted in the first equation? 2x – 2y = –12 4x – y = 3 a. –4x – 3 b. –4x + 3 c. 4x – 3 <--- i chose d. 4x + 3

OpenStudy (anonymous):

3. Use the elimination method to solve the following system of equations. 2x + 4y – 3z = –7 3x + y + 4z = –12 x + 3y + 4z = 4 a. (–6, –2, –1) <---- i chose b. (6, –2, –1) c. (–6, 2, –1) d. (–6, 2, 1) 4. Use the substitution method to solve the following system of equations. x + 2y – z = 7 4x – y + 3z = –2 2x + 2y – z = 9 a. (2, 1, 3) b. (2, 1, –3) <---- i chose c. (–2, 1, –3) d. (–2, –1, 3)

OpenStudy (ash2326):

questions 1 2 and 4 are correct Check the ques 3 again @bombshellbri

OpenStudy (anonymous):

thank you. and i tried it again, would it be -6, 2, 1? @ash2326

OpenStudy (ash2326):

correct:D

OpenStudy (anonymous):

Yes! Thanks again (: